# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v5

@@ I Introduction
The notable progress in generative models, such as GANs [ffhq] and Diffusion models [adm], is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
@@
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [wang2020cnn, zhong2023rich] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging. Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [I-FGSM, luo2022frequency], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts. This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints. Consequently, transferring adversarial patterns across such diverse representations is far less straightforward. Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [carlini2020evading, hussain2021adversarial, neekhara2021adversarial, hou2023evading, jia2022exploring]. These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces. In contrast, AIGI detectors are designed to identify any AI-generated content produced by a wide range of generative techniques, including GANs, diffusion models, and autoregressive models. The content extends beyond human faces to encompass non-human entities, landscapes, abstract art, and more. These differences make designing attacks for AIGI detection more challenging than Deepfake detection.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [dzanic2020fourier, frank2020leveraging] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detecting AI-generated images[zhu2024genimage], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors, e.g., transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
@@ II Related Work
AI-Generated Image Detection: The rapid progress of generative models, from early GANs [ffhq] to recent diffusion models [adm], has raised substantial risks of disinformation. While early work focused primarily on detecting fake human faces [miao2025multi, miao2025mixture], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection. AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images. Data-driven methods [wang2020cnn, feng2025deepfake] achieve strong in-distribution performance but generalize poorly to unseen generators. To address this, frequency-domain approaches [miao2022hierarchical, miao2023f] exploit forgery traces, while spatial-domain methods learn local forgery features [guo2023ldfnet]. Others leverage fingerprint representations from noise patterns [liu2022detecting]. Parameter-efficient strategies based on frozen pre-trained models have also been explored [tan2023learning], along with approaches exploiting diffusion reconstruction error [ricker2024aeroblade], teacher–student discrepancies [zhu2023gendet], or diffusion noise [zhang2023diffusion].
Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [ran2024adaptive, I-FGSM], multimedia communication [gao2024deepspoof], computational imaging [liang2025understanding] and human activity recognition [diao2024understanding, diaotasar, diao2021basar]. Gradient-based attack methods [I-FGSM, pgd] exploit gradients to mislead models. To improve adversarial transferability, frequency-domain approaches [luo2022frequency] perturb spectral representations, while ensemble-based attacks [dong2018boosting, svre] aggregate gradients or perturbations from multiple models. Very recently, their impact on Deepfake detection has been explored. Gradient sign-based attacks [hussain2021adversarial, I-FGSM, fgsm], latent space manipulations [li2021exploring], and black-box evaluations [carlini2020evading, neekhara2021adversarial] reveal vulnerabilities in existing detectors. Other works leverage frequency-domain perturbations [jia2022exploring], natural degradation noise [hou2023evading], or fingerprint removal strategies such as FakePolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader]. Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [zhou2024stealthdiffusion]. In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
@@ III-A Preliminaries
where $L$ is the binary cross-entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD [pgd] or I-FGSM [I-FGSM]:
@@ III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [wang2020cnn, zhong2023rich]. While these subtle clues are invisible in the spatial domain, a series of works [frank2020leveraging] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement the discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [long2022frequency] to visualize the sensitive components of real and fake images across different models:
@@
where $\Gamma(\cdot)$ denotes the random spectrum transformation [long2022frequency]. $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attacks in the frequency domain via spectrum transformation:
@@ III-C Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 5 can simulate different substitute models with a homogeneous architecture [long2022frequency], it shows limited transferability when applied to heterogeneous architectures, e.g., transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e., exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 5 by minimizing the Bayesian posterior predictive distribution:
@@ III-C1 Post-train Bayesian Strategy
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high-dimensional distribution due to a very large number of parameters of DNNs [izmailov2021bayesian]. Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximately sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high-dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [zhu2024genimage, he2021forgerynet]. From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
@@ III-C2 Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[springenberg_bayesian_2016] to sample appended model $\theta^{\prime}$ in each iteration:
@@
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [springenberg_bayesian_2016].
@@ III-D Hybrid Adversarial Attack
Despite detecting fake fingerprints in the frequency domain, some works also extract fingerprint features in the spatial domain [wang2020cnn]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
@@ IV-A Experimental Settings
Datasets: We chose three generated image datasets created by a wide range of generative models. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [wang2020cnn], containing 360k real images from LSUN and 360k fake images generated by ProGAN [progan]. GenImage [zhu2024genimage] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [zhu2024genimage], we employ a subset of GenImage, collecting 162k real images from Imagenet [deng2009imagenet] and 162k Stable Diffusion(SD) V1.4 [sd] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [shamshad2023evading]. The Deepfake dataset consists of 50k real face images from FFHQ [ffhq] and 50k generated face images generated by StyleGAN2 [ffhq]. After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [wang2020cnn], MobileNet [howard2017mobilenets], EfficientNet [tan2019efficientnet] and DenseNet [huang2017densely], and ViT-based detectors Vision Transformer(ViT) [vit] and Swin-Transformer(Swin-ViT) [liu2021swin]. For evaluating on various detection methods, we use frequency-based detectors DCTA [frank2020leveraging], Spec [zhang2019detecting], FreqNet [tan2024frequency], and FreqMask [doloriel2024frequency], gradient-based detectors LGrad [tan2023learning], CLIP-based detectors UnivFD [ojha2023towards], and diffusion-based detectors DNF [zhang2023diffusion], DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade]. In addition, LNP [liu2022detecting] extracts the noise pattern of images and GramNet [GramNet] learns the global texture representation. We also consider them as victim models. The datasets employed for training each detector are summarized in Table I. Note that AEROBLADE is a training-free method and not trained on the listed datasets.
@@
Compared Methods: We adopt gradient-based methods, I-FGSM [I-FGSM], PGD [pgd], MI [dong2018boosting] and frequency-based methods S2[long2022frequency], SSAH [luo2022frequency] and ensemble-based methods ENS [dong2018boosting] and SVRE [svre]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [jia2022exploring], Fakepolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
@@
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [wang2020cnn]. Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [wang2020cnn] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [wang2020cnn]. On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [bbc]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [long2022frequency]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.

### v6

@@ I Introduction
The notable progress in generative models, such as GANs [27] and Diffusion models [5], is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
@@
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [54, 63] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging. Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [28, 35], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts. This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints. Consequently, transferring adversarial patterns across such diverse representations is far less straightforward. Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [3, 23, 42, 19, 25]. These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces. In contrast, AIGI detectors are designed to identify any AI-generated content produced by a wide range of generative techniques, including GANs, diffusion models, and autoregressive models. The content extends beyond human faces to encompass non-human entities, landscapes, abstract art, and more. These differences make designing attacks for AIGI detection more challenging than Deepfake detection.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [12, 14] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detecting AI-generated images[66], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors, e.g., transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
@@ II Related Work
AI-Generated Image Detection: The rapid progress of generative models, from early GANs [27] to recent diffusion models [5], has raised substantial risks of disinformation. While early work focused primarily on detecting fake human faces [38, 37], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection. AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images. Data-driven methods [54, 13] achieve strong in-distribution performance but generalize poorly to unseen generators. To address this, frequency-domain approaches [40, 39] exploit forgery traces, while spatial-domain methods learn local forgery features [17]. Others leverage fingerprint representations from noise patterns [31]. Parameter-efficient strategies based on frozen pre-trained models have also been explored [51], along with approaches exploiting diffusion reconstruction error [46], teacher–student discrepancies [65], or diffusion noise [62].
Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [45, 28], multimedia communication [15], computational imaging [30] and human activity recognition [7, 8, 6]. Gradient-based attack methods [28, 36] exploit gradients to mislead models. To improve adversarial transferability, frequency-domain approaches [35] perturb spectral representations, while ensemble-based attacks [10, 58] aggregate gradients or perturbations from multiple models. Very recently, their impact on Deepfake detection has been explored. Gradient sign-based attacks [23, 28, 16], latent space manipulations [29], and black-box evaluations [3, 42] reveal vulnerabilities in existing detectors. Other works leverage frequency-domain perturbations [25], natural degradation noise [19], or fingerprint removal strategies such as FakePolisher [22] and TraceEvader [56]. Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [64]. In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
@@ III-A Preliminaries
where $L$ is the binary cross-entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD [36] or I-FGSM [28]:
@@ III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [54, 63]. While these subtle clues are invisible in the spatial domain, a series of works [14] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement the discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [34] to visualize the sensitive components of real and fake images across different models:
@@
where $\Gamma(\cdot)$ denotes the random spectrum transformation [34]. $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attacks in the frequency domain via spectrum transformation:
@@ III-C Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 5 can simulate different substitute models with a homogeneous architecture [34], it shows limited transferability when applied to heterogeneous architectures, e.g., transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e., exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 5 by minimizing the Bayesian posterior predictive distribution:
@@ III-C1 Post-train Bayesian Strategy
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high-dimensional distribution due to a very large number of parameters of DNNs [24]. Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximately sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high-dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [66, 18]. From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
@@ III-C2 Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[49] to sample appended model $\theta^{\prime}$ in each iteration:
@@
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [49].
@@ III-D Hybrid Adversarial Attack
Despite detecting fake fingerprints in the frequency domain, some works also extract fingerprint features in the spatial domain [54]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
@@ IV-A Experimental Settings
Datasets: We chose three generated image datasets created by a wide range of generative models. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [54], containing 360k real images from LSUN and 360k fake images generated by ProGAN [26]. GenImage [66] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [66], we employ a subset of GenImage, collecting 162k real images from Imagenet [4] and 162k Stable Diffusion(SD) V1.4 [47] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [48]. The Deepfake dataset consists of 50k real face images from FFHQ [27] and 50k generated face images generated by StyleGAN2 [27]. After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [54], MobileNet [20], EfficientNet [52] and DenseNet [21], and ViT-based detectors Vision Transformer(ViT) [11] and Swin-Transformer(Swin-ViT) [32]. For evaluating on various detection methods, we use frequency-based detectors DCTA [14], Spec [61], FreqNet [50], and FreqMask [9], gradient-based detectors LGrad [51], CLIP-based detectors UnivFD [43], and diffusion-based detectors DNF [62], DIRE [55] and AEROBLADE [46]. In addition, LNP [31] extracts the noise pattern of images and GramNet [33] learns the global texture representation. We also consider them as victim models. The datasets employed for training each detector are summarized in Table I. Note that AEROBLADE is a training-free method and not trained on the listed datasets.
@@
Compared Methods: We adopt gradient-based methods, I-FGSM [28], PGD [36], MI [10] and frequency-based methods S2[34], SSAH [35] and ensemble-based methods ENS [10] and SVRE [58]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [25], Fakepolisher [22] and TraceEvader [56] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
@@
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [54]. Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [54] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [54]. On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [53]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [34]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.